OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!! factor and check a^7b^7+1

OpenStudy (anonymous):

no answer choice

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

here answer b = {-7, -5}

OpenStudy (anonymous):

@-AngleWarrior-

OpenStudy (jdoe0001):

\(\large a^7b^7+1?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i gave u the answer

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

or hm gimme a sec

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

1 + a7b7

OpenStudy (jdoe0001):

\(\large {\textit{if "n" is an odd value then} \\ \quad \\ a^{\color{brown}{ n}}+b^{\color{brown}{ n}}=(a+b)(a^{{\color{brown}{ n}}-1}-a^{{\color{brown}{ n}}-2}b^1+a^{{\color{brown}{ n}}-3}b^2\quad ...\quad -a^1b^{{\color{brown}{ n}}-2}+b^{{\color{brown}{ n}}-1})}\) does that make any sense to you? I assume you may have covered that, or your book should have it

OpenStudy (jdoe0001):

well, \(\textit{if "n" is an odd value then} \\ \quad \\ a^{\color{brown}{ n}}+b^{\color{brown}{ n}}=(a+b)(a^{{\color{brown}{ n}}-1}-a^{{\color{brown}{ n}}-2}b^1+a^{{\color{brown}{ n}}-3}b^2\quad ...\quad -a^1b^{{\color{brown}{ n}}-2}+b^{{\color{brown}{ n}}-1})\)

OpenStudy (anonymous):

Yeah it makes since.

OpenStudy (jdoe0001):

keeping in mind that \(\bf a^7b^7+1\implies (ab)^7+1\implies (ab)^7+1^7\) recall that \(\bf 1^7 = 1\) thus

OpenStudy (jdoe0001):

so.. there :), just factor that, and that'd give you 2 factors then